a train leaves Havre at 5:30 and heads west at 40/mph. Half an hour later another train leaves Havre heading east at 60/mph what time will it be when the two trains are 200 miles apart
Answer by checkley71(8403)

40T+60(T-.5)=200
40T+60T-30=200
100T=200+30
100T=230
T=230/100
T=2.3 OR 2 HOURS & 18 MINUTES (5:30+2:18=7:48) THEY WILL BE 200 MILES APART.
PROOF
40*2.3+60(2.3-.5)=200
92+60*1.8=200
92+108=200
200=200